# A phase transition for the two-dimensional random field Ising/FK-Ising model

### v1

@@ Appendix A A Some LDP results for FK-Ising model at criticality
The proof of Theorem A.1 follows the framework of [24].
@@
Given the event $C(l,2l)$ happens, there is an open path $\gamma$ in $\mathcal{A}(l,2l)$ separating its inner and outer boundaries, thus $\gamma$ also separates $\partial_{ext}\Lambda_{k}$ and $\partial_{int}\Lambda_{m}.$ As a result, if both $D(k,2l)$ and $D(l,m)$ happen, this two crossing should be connected together by $\gamma.$ (See Figure 3 for illustration) Combined with CBC, we have
@@
Finally, by the RSW theory (see [17, Theorem 1.1]), we know there is a positive constant $c^{\prime}>0$ such that
@@
$\mathcal{V}_{n}:=\left\{v\in\Lambda_{n}\,|\,v\leftrightarrow\partial_{int}\Lambda_{2n}\right\}$
to be the set of vertices in $\Lambda_{n}$ which are connected to $\partial_{int}\Lambda_{2n}$. Then [24, Theorem 1.5] and (A.1) show
@@
PART 1:
In the first part, we show that all the clusters except for the boundary cluster cannot be too large, the proof is really similar to the proof of [5, Proposition 6.3]. In order to be self-contained, we give a short proof here. Without loss of generality, we may assume that $N=2^{n}$ and define
$\mathscr{B}_{l}:=\{\Lambda_{2^{l-1}}(u):u\in(2^{l}\mathbb{Z})^{2}\}$
to be a box partition of $\Lambda_{N}$ with side length $2^{l}.$ Then we write
$\mathscr{B}:=\bigcup_{l=1}^{n}\mathscr{B}_{l}.$
Now for a box $\Lambda_{2^{l-1}(u)}\in\mathscr{B}_{l},$ we write $L=2^{l-1}$ for short and define
$\mathcal{V}_{L}(u):=\{v\in\Lambda_{2^{l-1}(u)}\mid v\leftrightarrow\Lambda_{2^{l}(u)}\}.$
We will call the box $\Lambda_{2^{l-1}(u)}$ is crowded if
$|\mathcal{V}_{L}(u)|\geq\frac{1}{289}N^{\frac{15}{8}}x$
and we have the following claim:
@@
$\sum_{l=2}^{l_{0}}C_{1}\exp\left(-\frac{1}{C_{1}}N^{\frac{15}{8}}x\right)\cdot(\frac{N}{L})^{2}+\sum_{l=l_{0}+1}^{\lfloor\log_{2}N\rfloor}C_{1}\exp\left(-\frac{N^{30}x^{16}}{C_{1}L^{30}}\right)\cdot(\frac{N}{L})^{2}\leq C_{2}\exp\left(-C_{2}^{-1}x^{16}\right).$
PART 2:
@@
For convenience, we use $\mathcal{T}_{l}(l=7k+a-1,\cdots,8k-2)$ to denote the union of boxes in $\mathscr{T}_{l}.$ Since the number of points not contained in $\cup_{l=7k+a-1}^{8k-2}\mathcal{T}_{l}$ is $N^{\frac{15}{8}}2^{a-1}<\frac{N{\frac{15}{8}}x}{2}$, we only need to consider the intersection of $\mathcal{C}_{*}$ and $\cup_{l=7k+a-1}^{8k-2}\mathcal{T}_{l}$. Let $C_{3}$ be an absolute constant such that $\sum_{l=7k+a-1}^{8k-2}C_{3}2^{\frac{-8k+l}{2}}\leq 1$. Then we have
@@
For a configuration $\omega\in\Xi$, define
$\mathscr{M}(\omega):=\sum_{\mathcal{C}\in\mathfrak{C}(\omega)}|\mathcal{C}|^{2},$
to be the summation of the squares of the number of the clusters in $\omega.$
@@
where $k_{0}=\max\{k\in\mathbb{N}~{}|~{}2^{k}<N^{\frac{15}{8}}\sqrt{x}\}$. As in the proof of Theorem A.1, the factor $2^{\frac{k-k_{0}}{2}}$ is introduced to ensure the sequence is summable. We will first consider a dividing procedure of $\Lambda_{N}$. To be precise, we fix an integer $M=M(k)>0$ and consider the following horizontal and vertical line segments inside $\Lambda_{N}$:
@@
$\Psi\left(\{\sum|\mathcal{C}^{*}_{(i,j)}|\geq C_{0}\frac{VN^{2}}{M^{2}}\right\}\cap A^{*})\leq\Phi\left(\text{at least }C_{0}\frac{VN^{2}}{2^{k}M^{2}}~{}\mathcal{N}(i,j)\text{ happens}\right).$ (A.19)
@@
$\Psi\left(\{\sum|\mathcal{C}^{*}_{(i,j)}|\geq C_{0}\frac{VN^{2}}{M^{2}}\}\cap A^{*}\right)$
@@
The proof of Theorem A.6 is the same as the proof of [24, Theorem 1.4].
@@ Appendix B B Some analytic results
$\mathsf{var}(X_{K})$ $\geq\mathsf{var}(X)-{\mathbb{E}}~{}(X^{1}-X^{2})^{2}(\mathbf{1}_{|X^{1}|\geq K\text{ or }|X^{2}|\geq K})$
$=\mathsf{var}(X)-2{\mathbb{E}}(X^{1}-X^{2})^{2}\mathbf{1}_{|X^{1}|\geq K,|X^{2}|<K}-{\mathbb{E}}(X^{1}-X^{2})^{2}\mathbf{1}_{|X^{1}|\geq K,|X^{2}|\geq K}$
$\geq\mathsf{var}(X)-2{\mathbb{E}}(|X^{1}|+K)^{2}\mathbf{1}_{|X^{1}|\geq K}-2{\mathbb{E}}[(X^{1})^{2}+(X^{2})^{2}]\mathbf{1}_{|X^{1}|\geq K,|X^{2}|\geq K}$
$\geq\mathsf{var}(X)-2{\mathbb{E}}(2|X^{1}|)^{2}\mathbf{1}_{|X^{1}|\geq K}-4{\mathbb{E}}[(X^{1})^{2}]\mathbf{1}_{|X^{1}|\geq K}$
@@
$=$ ${\mathbb{E}}(X_{K}(\omega_{1})-X_{K}(\omega_{2}))^{2}\mathbf{1}_{\omega_{1}\in\Omega_{1}\cup\Omega_{2}\text{ or }\omega_{2}\in\Omega_{1}\cup\Omega_{2}}+{\mathbb{E}}(X_{K}(\omega_{1})-X_{K}(\omega_{2}))^{2}\mathbf{1}_{\omega_{1},\omega_{2}\in\Omega_{1}^{c}\cap\Omega_{2}^{c}}$
@@
Combined with the fact that $\mathsf{var}(X_{K})\geq\frac{c_{1}}{2}$ ,it yields that
$|\min_{\omega_{1}\in\Omega_{1}}X_{K}(\omega_{1})-\max_{\omega_{2}\in\Omega_{2}}X_{K}(\omega_{2})|\geq\sqrt{\frac{c_{1}}{4}}.$
Thus, we have for any $\omega_{1}\in\Omega_{1},\omega_{2}\in\Omega_{2}$ that
@@
$\big{|}X(\omega)-\log\left({\mathbb{E}}~{}e^{X}\right)\big{|}\geq\frac{\sqrt{c_{1}}}{4}.$
@@
For any constants $c_{1},c_{2}>0$, there exists a constant $c_{3}=c_{3}(c_{1},c_{2})>0$ such that for any random variable $X$ with ${\mathbb{E}}\exp(X)\leq c_{1}$ and $\mathsf{var}(X)\geq c_{2}$, we have $\log\big{(}{\mathbb{E}}\exp(X)\big{)}\geq{\mathbb{E}}X+c_{3}$.
@@
The desired result thus holds since ${\mathbb{E}}\exp(X)\leq c_{1}$ and $\mathsf{var}(X)\geq c_{2}$. ∎
@@
$\mathbb{P}\left(\Bigg{|}\sum_{i=1}^{n}a_{i}Z_{i}\Bigg{|}\geq x\right)\leq c\exp\left(-\frac{x^{\frac{1}{4}}}{c(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{8}}}\right).$
@@
The proof of (B.3) is similar to that of [11, Lemma B.1], so we omit it here. The proof of (B.4) comes from the fact that $X_{i}^{4}=3+6Y_{i}+Z_{i}$ and combining with (B.3). ∎
@@
Let $X_{1},X_{2},\cdots,X_{n}$ be independent random variables with a uniform exponential moment bound, i.e., ${\mathbb{E}}\exp(|X_{i}|)\leq C_{1}$ for $i\in[1,n]$. Then for any $c_{2}>0$, there exists a constant $c_{3}>0$ such that
$\mathbb{P}\left(\Bigg{|}\sum_{i=1}^{n}X_{i}-{\mathbb{E}}X_{i}\Bigg{|}\geq c_{2}n\right)\leq\exp(-c_{3}n).$
@@ Appendix C C Some results on FK-Ising model without disorder
$M$ $=\sum_{x\in\Omega}\mathbf{1}_{\{x\stackrel{{\scriptstyle\omega}}{{\longleftrightarrow}}\partial_{int}\Omega\}}-\sum_{x}\mathbf{1}_{\{x\stackrel{{\scriptstyle\omega}}{{\longleftrightarrow}}\partial_{int}\Omega\}}\mathbf{1}_{\{x\stackrel{{\scriptstyle\tilde{\omega}}}{{\longleftrightarrow}}\partial_{int}\Omega\}}$
@@ Acknowledgements
We thank Jian Ding for several insightful discussions. We thank Xinyi Li and João Maia for reading an early version of this paper and for providing many valuable suggestions. F. Huang is supported by the Beijing Natural Science Foundation Undergraduate Initiating Research Program (Grant No. QY23006). All authors are partially supported by the National Natural Science Foundation of China (NSFC) through the Tianyuan Mathematics Key Program (Grant No. 12226001) and the NSFC Key Program (Grant No. 12231002).

### v2

@@ Appendix A A Some LDP results for FK-Ising model at criticality
As we remarked above, a continuous version of Theorem A.1 has been shown in [11], and their proof cannot be applied to the discrete settings. So we follow the framework of [28] to prove Theorem A.1.
@@
Given the event $C(l,2l)$ happens, there is an open path $\gamma$ in $\mathcal{A}(l,2l)$ separating its inner and outer boundaries, thus $\gamma$ also separates $\partial_{ext}\Lambda_{k}$ and $\partial_{int}\Lambda_{m}.$ As a result, if both $D(k,2l)$ and $D(l,m)$ happen, this two crossing should be connected together by $\gamma$ (See Figure 4 for illustration). Combined with CBC, we have
@@
Finally, by the RSW theory (see [21, Theorem 1.1]), we know there is a positive constant $c^{\prime}>0$ such that
@@
$\mathcal{V}_{n}:=\left\{v\in\Lambda_{n}\,:\,v\leftrightarrow\partial_{int}\Lambda_{2n}\right\}$
to be the set of vertices in $\Lambda_{n}$ which are connected to $\partial_{int}\Lambda_{2n}$. Then [28, Theorem 1.5] and (A.1) show
@@
PART 1: Controlling the bulk cluster
In the first part, we show that all the clusters except for the boundary cluster cannot be too large, the proof is really similar to the proof of [5, Proposition 6.3]. In order to be self-contained, we give a short proof here.
Without loss of generality, we may assume that $N=2^{n}$ and define $\mathscr{B}_{l}:=\{\Lambda_{2^{l-1}}(u):u\in(2^{l}\mathbb{Z})^{2}\}$ to be a box partition of $\Lambda_{N}$ with side length $2^{l}.$ Then we write $\mathscr{B}:=\bigcup_{l=1}^{n}\mathscr{B}_{l}.$ Now for a box $\Lambda_{2^{l-1}(u)}\in\mathscr{B}_{l},$ we write $L=2^{l-1}$ for short and define $\mathcal{V}_{L}(u):=\{v\in\Lambda_{2^{l-1}(u)}:v\leftrightarrow\Lambda_{2^{l}(u)}\}.$ We will call the box $\Lambda_{2^{l-1}(u)}$ is crowded if $|\mathcal{V}_{L}(u)|\geq\frac{1}{289}N^{\frac{15}{8}}x$ and we have the following claim:
@@
$\sum_{l=2}^{l_{0}}C_{1}\exp\left(-\frac{1}{C_{1}}N^{\frac{15}{8}}x\right)\cdot\left(\frac{N}{L}\right)^{2}+\sum_{l=l_{0}+1}^{\lfloor\log_{2}N\rfloor}C_{1}\exp\left(-\frac{N^{30}x^{16}}{C_{1}L^{30}}\right)\cdot\left(\frac{N}{L}\right)^{2}\leq C_{2}\exp\left(-C_{2}^{-1}x^{16}\right).$
PART 2: Controlling the boundary cluster
@@
For convenience, we use $\mathcal{T}_{l}(l=7k+a-1,\cdots,8k-2)$ to denote the union of boxes in $\mathscr{T}_{l}.$ Since the number of points not contained in $\cup_{l=7k+a-1}^{8k-2}\mathcal{T}_{l}$ is $N^{\frac{15}{8}}2^{a-1}<\frac{N^{\frac{15}{8}}x}{2}$, we only need to consider the intersection of $\mathcal{C}_{*}$ and $\cup_{l=7k+a-1}^{8k-2}\mathcal{T}_{l}$. Let $C_{3}$ be an absolute constant such that $\sum_{l=7k+a-1}^{8k-2}C_{3}2^{\frac{-8k+l}{2}}\leq 1$. Then we have
@@
For a configuration $\omega\in\Xi$, define $\mathscr{M}(\omega):=\sum_{\mathcal{C}\in\mathfrak{C}(\omega)}|\mathcal{C}|^{2},$ to be the summation of the squares of the number of clusters in $\omega.$
@@
where $k_{0}=\max\{k\in\mathbb{N}:2^{k}<N^{\frac{15}{8}}\sqrt{x}\}$. As in the proof of Theorem A.1, the factor $2^{\frac{k-k_{0}}{2}}$ is introduced to ensure the sequence is summable. We will first consider a dividing procedure of $\Lambda_{N}$. To be precise, we fix an integer $M=M(k)>0$ and consider the following horizontal and vertical line segments inside $\Lambda_{N}$:
@@
$\Psi\left(\left\{\sum|\mathcal{C}^{*}_{(i,j)}|\geq C_{0}\frac{VN^{2}}{M^{2}}\right\}\cap A^{*}\right)\leq\Phi\left(\text{at least }C_{0}\frac{VN^{2}}{2^{k}M^{2}}~{}\mathcal{N}(i,j)\text{ happens}\right).$ (A.19)
@@
$\Psi\left(\left\{\sum|\mathcal{C}^{*}_{(i,j)}|\geq C_{0}\frac{VN^{2}}{M^{2}}\right\}\cap A^{*}\right)$
@@
The proof of Theorem A.6 is the same as the proof of [28, Theorem 1.4].
@@ Appendix B B Some analytic results
$\mathsf{var}(X_{K})$ $\geq\mathsf{var}(X)-{\mathbb{E}}~{}(X^{1}-X^{2})^{2}\mathbf{1}_{\{|X^{1}|\geq K\text{ or }|X^{2}|\geq K\}}$
$=\mathsf{var}(X)-2{\mathbb{E}}(X^{1}-X^{2})^{2}\mathbf{1}_{\{|X^{1}|\geq K,|X^{2}|<K\}}-{\mathbb{E}}(X^{1}-X^{2})^{2}\mathbf{1}_{\{|X^{1}|\geq K,|X^{2}|\geq K\}}$
$\geq\mathsf{var}(X)-2{\mathbb{E}}(|X^{1}|+K)^{2}\mathbf{1}_{\{|X^{1}|\geq K\}}-2{\mathbb{E}}[(X^{1})^{2}+(X^{2})^{2}]\mathbf{1}_{\{|X^{1}|\geq K,|X^{2}|\geq K\}}$
$\geq\mathsf{var}(X)-2{\mathbb{E}}(2|X^{1}|)^{2}\mathbf{1}_{\{|X^{1}|\geq K\}}-4{\mathbb{E}}[(X^{1})^{2}]\mathbf{1}_{\{|X^{1}|\geq K\}}$
@@
$=$ ${\mathbb{E}}(X_{K}(\omega_{1})-X_{K}(\omega_{2}))^{2}\mathbf{1}_{\{\omega_{1}\in\Omega_{1}\cup\Omega_{2}\text{ or }\omega_{2}\in\Omega_{1}\cup\Omega_{2}\}}+{\mathbb{E}}(X_{K}(\omega_{1})-X_{K}(\omega_{2}))^{2}\mathbf{1}_{\{\omega_{1},\omega_{2}\in\Omega_{1}^{c}\cap\Omega_{2}^{c}\}}$
@@
Combined with the fact that $\mathsf{var}(X_{K})\geq\frac{c_{1}}{2}$ ,it yields that $\big{|}\min_{\omega_{1}\in\Omega_{1}}X_{K}(\omega_{1})-\max_{\omega_{2}\in\Omega_{2}}X_{K}(\omega_{2})\big{|}\\ \geq\sqrt{\frac{c_{1}}{4}}.$Thus, we have for any $\omega_{1}\in\Omega_{1},\omega_{2}\in\Omega_{2}$ that
@@
$\big{|}X(\omega)-\ln\left({\mathbb{E}}~{}e^{X}\right)\big{|}\geq\frac{\sqrt{c_{1}}}{4}.$
@@
For any constant $c_{1}>0$, there exists a constant $c_{2}>0$ only relies on $c_{1}$ such that for any random variable $X$ with ${\mathbb{E}}\exp(X)<\infty$ and $\mathsf{var}(X)\geq c_{1}$, we have $\ln\big{(}{\mathbb{E}}\exp(X)\big{)}\geq{\mathbb{E}}X+c_{2}$.
@@
The desired result thus holds since $\mathsf{var}(X)\geq c_{2}$. ∎
@@
$\mathbb{P}\left(\Bigg{|}\sum_{i=1}^{n}a_{i}Z_{i}\Bigg{|}\geq x\right)\leq c\exp\left(-\frac{x^{\frac{1}{4}}}{c(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{8}}}\right).$ (B.4)
@@
The proofs of (B.3) and (B.4) are similar to that of [15, Lemma B.1] (see [31, Lemma 1] for an alternative proof to the first inequality), so we only give a sketch for (B.4) here. We start from controlling the moments of $\Big{|}\sum_{i=1}^{n}a_{i}Z_{i}\Big{|}$. By hypercontractivity for Guassian variables [33] (see also [34, Theorem 1.4.1]), we get that
${\mathbb{E}}\Big{|}\sum_{i=1}^{n}a_{i}Z_{i}\Big{|}^{p}\leq(p-1)^{2p}\Big{(}{\mathbb{E}}(\sum_{i=1}^{n}a_{i}Z_{i})^{2}\Big{)}^{\frac{p}{2}}=(p-1)^{2p}\Big{(}\sum_{i=1}^{n}a_{i}^{2}\Big{)}^{\frac{p}{2}}.$ (B.6)
By [15, Equation (B.8)], we can control the exponential moment of $\Big{|}\sum_{i=1}^{n}a_{i}Z_{i}\Big{|}^{\frac{1}{4}}$
$\exp(t\Big{|}\sum_{i=1}^{n}a_{i}Z_{i}\Big{|}^{\frac{1}{4}})=\sum_{p=0}^{\infty}{\mathbb{E}}\Big{|}\sum_{i=1}^{n}a_{i}Z_{i}\Big{|}^{\frac{p}{4}}\cdot t^{p}/p!$ $\leq\sum_{p=0}^{\infty}(\sum_{i=1}^{n}a_{i}^{2})^{\frac{p}{8}}(t)^{p}$
$=\frac{1}{1-(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{8}}\cdot t}$
where $t\in(0,\frac{1}{(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{8}}})$. Choosing $t=\frac{1}{2(\sum_{i=1}^{n}a_{i}^{2})^{\frac{1}{8}}}$ and applying an exponential version of the Markov inequality completes the proof of (B.4). The proof of (B.5) comes from the fact that $X_{i}^{4}=3+6Y_{i}+Z_{i}$ and combining with (B.3) and (B.4). ∎
@@
Let $X_{1},X_{2},\cdots,X_{n}$ be independent random variables with a uniform exponential moment bound, i.e., ${\mathbb{E}}\exp(|X_{i}|)\leq C_{1}$ for $1\leq i\leq n$. Then for any $x>0$, there exists a constant $c>0$ relies on $C_{1}$ but not $x$ such that
$\mathbb{P}\left(\Bigg{|}\sum_{i=1}^{n}X_{i}-{\mathbb{E}}X_{i}\Bigg{|}\geq x\right)\leq\exp(-\min\{\frac{x^{2}}{nc},\frac{x}{c}\}).$
@@ Acknowledgements
We thank Jian Ding for several insightful discussions. We thank Xinyi Li and João Maia for reading an early version of this paper and for providing many valuable suggestions. We thank Federico Camia and Jianping Jiang for reminding us the continuous version of Theorem A.1. F. Huang is supported by the Beijing Natural Science Foundation Undergraduate Initiating Research Program (Grant No. QY23006). All authors are partially supported by the National Natural Science Foundation of China (NSFC) through the Tianyuan Mathematics Key Program (Grant No. 12226001) and the NSFC Key Program (Grant No. 12231002).